# Embodied AI in Urban Spaces

### v1

@@ Abstract
Public urban spaces like streetscapes and plazas serve residents and accommodate social life in all its vibrant variations. Recent advances in Robotics and Embodied AI make public urban spaces no longer exclusive to humans. Food delivery bots and electric wheelchairs have started sharing sidewalks with pedestrians, while diverse robot dogs and humanoids have recently emerged in the street. Ensuring the generalizability and safety of these forthcoming mobile machines is crucial when navigating through the bustling streets in urban spaces. In this work, we present MetaUrban, a compositional simulation platform for Embodied AI research in urban spaces. MetaUrban can construct an infinite number of interactive urban scenes from compositional elements, covering a vast array of ground plans, object placements, pedestrians, vulnerable road users, and other mobile agents’ appearances and dynamics. We design point navigation and social navigation tasks as the pilot study using MetaUrban for embodied AI research and establish various baselines of Reinforcement Learning and Imitation Learning. Experiments demonstrate that the compositional nature of the simulated environments can substantially improve the generalizability and safety of the trained mobile agents. MetaUrban will be made publicly available to provide more research opportunities and foster safe and trustworthy embodied AI in urban spaces.
@@ 1 Introduction
Public urban spaces vary widely in type, form, and size, encompassing streetscapes, plazas, and parks. They are crucial spaces for transit and transport, as well as providing opportunities to stage various social events. Since the early 20th century, the study of public urban spaces has long been a cornerstone of urban sociology [56, 32, 29] and planning [26, 78, 28]. For example, William H. Whyte, in his seminal work, “City - Rediscovering the Center” [79], revealed that the complexity and vibrant interaction in public spaces profoundly determine humans’ social life, underscoring the critical role these environments play in urban safety and vitality.
@@
Recent development of Robotics and Embodied AI makes the urban space no longer exclusive to humans. Various mobile machines have started emerging. For example, elders and physically disabled people maneuver electronic wheelchairs on the street, while food delivery bots navigate on the sidewalk to accomplish the last-mile food delivery task. Various mobile legged robots like robot dog Spot from Boston Dynamics and humanoid robot Optimus from Tesla are also forthcoming. We can thus imagine a future of public urban spaces that will be shared and co-habitated by humans and mobile machines driven by Embodied AI. Ensuring the generalizability and safety of these mobile machines becomes essential.
@@
Simulation platforms [35, 67, 70, 41, 73, 13, 18, 38, 44, 19] have played a crucial role in enabling systematic and scalable training of the embodied AI agents and the safety evaluation before real-world deployment. However, most of the existing simulators focus either on indoor household environments [61, 35, 67, 70, 41, 24] or outdoor driving environments [38, 44, 19]. For example, platforms like AI2-THOR [35], Habitat [67], and iGibson [70] are designed for household assistant robots in which the environments are mainly apartments or houses with furniture and appliances; platforms like SUMO [38], CARLA [19], and MetaDrive [44] are designed for research on autonomous driving and transportation. Yet, simulating urban spaces with diverse layouts and objects, complex dynamics of pedestrians, is much less explored.
@@
Distinct from the indoor household and driving environments, the urban space has unique characteristics. Let’s follow the adventure of a last-mile delivery bot, who aims to deliver a lunch order from a nearby pizzeria to the campus. First, it faces a long-horizon journey across several street blocks at a one-mile distance, with multifarious road hazards, such as fragmented curbs and rugged ground caused by tree roots on sidewalks. Then, it must safely navigate the cluttered street full of obstacles like trash bins, parked scooters, and potted plants. In addition, it needs to handle pedestrians and crowds properly to avoid collisions. It should also take special care around disabled people in wheelchairs. Thus, the layout diversity, object distribution, and dynamic complexity bring unique challenges to the design of simulation environments and the study of the generalizability and safety of Embodied AI agents operating in urban spaces.
@@
We present MetaUrban – a compositional simulation platform for Embodied AI research in urban spaces. First, we introduce Hierarchical Layout Generation, a procedural generation approach that can generate infinite layouts hierarchically from street blocks to sidewalks, functional zones, and object locations. It can generate scenes at an arbitrary scale with various connections and divisions of street blocks, object locations, and terrains, which are critical for improving the generalizability of trained agents. Then, we design the Scalable Object Retrieval, an automatic pipeline that can obtain an arbitrary number of high-quality objects with real-world distribution. We first compute the object category distribution from broad real-world data to form a description pool. Then, with the sampled descriptions from the pool, we effectively retrieve objects from large-scale 3D asset repositories with a VLM-based open-vocabulary searching schema. Finally, we propose the Cohabitant Populating method to generate complex dynamics in urban spaces. We first tailor recent 3D human and motion datasets to get 1,100 rigged pedestrian models, each with 2,314 movements. Then, to form safety-critical scenarios, we integrate Vulnerable Road Users (VRUs) like bikers, skateboarders, and scooter riders. To broaden the category of mobile machines in urban scenes, we include delivery bots, electric wheelchairs, robot dogs, and humanoid robots. Then, based on path planning algorithms, we can get complex trajectories among hundreds of environmental agents simultaneously with collision and deadlock avoidance. It is critical for enhancing the social conformity and safety of the mobile agents.
Based on MetaUrban, we construct a large-scale dataset, MetaUrban-12K, that includes 12,800 training scenes and 1,000 test scenes. We further create an unseen test set with 100 manually designed scenes to evaluate trained models’ generalizability. Besides, we provide 30,000 steps of high-quality expert demonstrations by humans and well-trained agents to enable offline reinforcement learning and imitation learning research. We design point navigation and social navigation tasks as a pilot study using MetaUrban for embodied AI research and establish extensive baselines for Reinforcement Learning, Safe Reinforcement Learning, Offline Reinforcement Learning, and Imitation Learning. Experiments demonstrate that the generalizability and safety of the trained mobile agents can be substantially improved by the compositional nature of the simulated environments. We will make the simulator available to enable more research opportunities for the community and foster safe and trustworthy embodied AI in urban spaces.
@@ 2 Related Work
Many simulation platforms have been developed for Embodied AI research, depending on the target environments – such as indoor homes and offices, driving freeways and roadways, and crowds in warehouses and squares. We compare representative ones with the proposed MetaUrban simulator.
@@
Platforms for indoor environments are mainly designed for household assistant robots, emphasizing the affordance, realism, and diversity of objects, as well as the interactivity of environments. VirtualHome [61] pivots towards simulating routine human activities at home. AI2-THOR [35] and its extensions, such as ManipulaTHOR [20], RoboTHOR [13], and ProcTHOR [18], focus on detailed agent-object interactions, dynamic object state changes, and procedural scene generation, alongside robust physics simulations. Habitat [67] offers environments reconstructed from 3D scans of real-world interiors. Its subsequent iterations, Habitat 2.0 [73] and Habitat 3.0 [63], introduce interactable objects and deformable humanoid agents, respectively. iGibson [70] provides photorealistic environments. Its upgrades, Gibson 2.0 [40], and OmniGibson [41], focus on household tasks with object state changes and a realistic physics simulation of everyday activities, respectively. ThreeDWorld [24] targets real-world physics by integrating high-fidelity simulations of liquids and deformable objects. However, unlike MetaUrban, these simulators are focused on indoor environments with particular tasks like object rearrangement and manipulation.
@@
Platforms for driving environments are mainly designed for autonomous vehicle research and development. Simulators like GTA V [50], Sim4CV [51], AIRSIM [69], CARLA [19], and its extension SUMMIT [7] offer realistic environments that mimic the physical world’s detailed visuals, weather conditions, and day-to-night transitions. Other simulators enhance efficiency and extensibility at the expense of visual realism, such as Udacity [75], DeepDrive [74], Highway-env [39], and DriverGym [37]. MetaDrive [44] trades off between visual quality and efficiency, offering a lightweight driving simulator that can support the research of generalizable RL algorithms for vehicles. Although some of the simulators [50, 19] involve traffic participants other than vehicles, such as pedestrians and cyclists, all of them focus on vehicle-centric driving scenarios and neglect environments and things happening in public urban spaces like sidewalks and plazas.
@@
Other than indoor and driving environments, social navigation platforms emphasize the social compatibility of robots. Simulators like Crowd-Nav [8], Gym-Collision-Avoidance [22], and Social-Gym 2.0 [71], model scenes and agents in 2D maps, focusing more on the development of path planning algorithms. Other simulators, such as HuNavSim [59], SEAN 2.0 [76], and SocNavBench [5], upgrade the environment to 3D space and introduce human pedestrians to support the development of more complex algorithms. However, social navigation platforms focus on crowd navigation, with oversimplified objects and surrounding environmental structures in the scenes. This proposed platform addresses the gap between the existing social navigation platforms and the real-world urban spaces regarding environmental diversity and complexity.
We will compare MetaUrban with other simulators below, through the scale, sensor, and feature dimensions. For the scale, MetaUrban can generate infinite scenes with a procedural generation pipeline. It also provides the largest number of humans (1,100) and movements (2,314) among all simulation environments. For the sensor, MetaUrban provides RGBD, semantic, and lidar, while acoustic is our next step to better support multi-model tasks. For the feature, different from other simulators, MetaUrban provides real-world distribution of the object’s categories and uses a more sophisticated path plan algorithm to get natural human trajectories. It also provides flexible user interfaces – mouse, keyboard and joystick and racing wheel, which vastly ease the collection of human expert demonstration data. MetaUrban uses PyBullet as its physical engine, which is open-source and highly accurate in physics simulation, providing a cost-effective and flexible solution for future developments. MetaUrban uses Panda3D [27] for rendering, which is a lightweight, open-source game engine with seamless Python integration, providing a flexible and accessible development environment. A detailed comparison table is included in the Appendix.
In summary, none of the recent simulation platforms have been constructed for urban spaces, and the proposed simulator differs from them significantly in terms of diverse layouts, objects, human dynamics, and different types of mobile agents like delivery robots, electric wheelchairs, robot dogs, humanoid robots, and vehicles, and their intricate interactions. We believe MetaUrban can provide a lot of new research opportunities for Embodied AI in urban settings.
@@ 3 MetaUrban Simulator
MetaUrban is a compositional simulation platform that can generate infinite training and evaluation environments for Embodied AI in urban spaces. Figure 2 depicts the procedural generation pipeline. MetaUrban uses a structured description script to create urban scenes. Based on the provided information about street blocks, sidewalks, objects, agents, and more, it starts with the street block map, then plans the ground layout by dividing different function zones, then places static objects, and finally populates dynamic agents.
@@
This section highlights three key designs in the MetaUrban simulator to support exhibiting three unique characteristics of urban spaces – diverse layouts, particular object distribution, and complex dynamics. Section 3.1 introduces Hierarchical Layout Generation, which can infinitely generate diverse layouts with different functional zone divisions and object locations that are critical for the generalizability of agents. Section 3.2 introduces Scalable Object Retrieval, which harnesses worldwide urban scene data to obtain real-world object distributions in different places, and then builds large-scale, high-quality static objects set with VLM-enabled open-vocabulary searching. It is useful for training agents specialized for urban scenes. Section 3.3 introduces Cohabitant Populating, in which we leverage the advancements in digital humans to enrich the appearances, movements, and trajectories of pedestrians and vulnerable road users, as well as incorporate other agents to form a vivid cohabiting environment. It is critical for improving the social conformity and safety of the mobile agents.
@@ 3.1 Hierarchical Layout Generation
The diversity of scene layout, i.e., the connection and categories of blocks, the specifications of sidewalks and crosswalks, as well as the placement of objects, is crucial for enhancing the generalizability of trained agents maneuvering in public spaces. In the hierarchical layout generation framework, we start by sampling the categories of street blocks and dividing sidewalks and crosswalks and then allocate various objects, with which we can get infinite urban scene layouts with arbitrary sizes and specifications of maps.
@@
As shown in Figure 3 (left), according to the Global Street Design Guide [31] provided by the Global Designing Cities Initiative, we divide the sidewalk into four functional zones – building zone, frontage zone, clear zone, and furnishing zone. Based on their different combinations of functional zones, we further construct 7 typical templates for sidewalks (Figure 3 (right)). To form a sidewalk, we can first sample the layout from the templates and then assign proportions for different function zones. For crosswalks, we provide candidates at the start and the end of each roadway, which support specifying the needed crosswalks or sampling them by a density parameter. Finally, roadways, sidewalks, and crosswalks can take a terrain map as substrate to form different ground situations.
@@
Object placement. After determining the ground layout, we can place objects on the ground. We divide objects into three classes. 1) Standard infrastructure, such as poles, trees, and signs, are placed periodically along the road. 2) Non-standard infrastructure, such as buildings, bonsai, and trash bins, are placed randomly in the designated function zones. 3) Clutter, such as drink cans, bags, and bicycles, are placed randomly across all functional zones. We can get different street styles by specifying an object pool while getting different compactness by specifying a density parameter. Figure 4 shows different objects placed with a sampled ground plan and objects’ location.
Figure 4: Object placement. Placing Different objects, with the same ground plan and objects’ location.
3.2 Scalable Object Retrieval

### v2

# Platform for Urban Micromobility

@@ Abstract
00footnotetext: * Equal contribution.
Public urban spaces like streetscapes and plazas serve residents and accommodate social life in all its vibrant variations. Recent advances in Robotics and Embodied AI make public urban spaces no longer exclusive to humans. Food delivery bots and electric wheelchairs have started sharing sidewalks with pedestrians, while robot dogs and humanoids have recently emerged in the street. Micromobility enabled by AI for short-distance travel in public urban spaces plays a crucial component in the future transportation system. Ensuring the generalizability and safety of AI models maneuvering mobile machines is essential. In this work, we present MetaUrban, a compositional simulation platform for the AI-driven urban micromobility research. MetaUrban can construct an infinite number of interactive urban scenes from compositional elements, covering a vast array of ground plans, object placements, pedestrians, vulnerable road users, and other mobile agents’ appearances and dynamics. We design point navigation and social navigation tasks as the pilot study using MetaUrban for urban micromobility research and establish various baselines of Reinforcement Learning and Imitation Learning. We conduct extensive evaluation across mobile machines, demonstrating that heterogeneous mechanical structures significantly influence the learning and execution of AI policies. We perform a thorough ablation study, showing that the compositional nature of the simulated environments can substantially improve the generalizability and safety of the trained mobile agents. MetaUrban will be made publicly available to provide research opportunities and foster safe and trustworthy embodied AI and micromobility in cities. The code and dataset will be released.
Figure 1: MetaUrban enables the construction of infinite interactive urban scenes, supports multiple sensors, and offers flexible user interfaces such as a mouse, keyboard, joystick, and racing wheel. The platform includes 10,000 diverse obstacles in urban scenes, 1,100 rigged human models each with 2,314 movements, vulnerable road users, mobile machines with varied mechanical structures, and a terrain generation system to create complex ground conditions. We highly recommend visiting our project page for video demonstrations.
@@ 1 Introduction
Public urban spaces (Whyte, 2012) vary widely in type, form, and size, encompassing streetscapes, plazas, and parks. They are crucial spaces for transit and transport (Geddes, 1949), as well as providing stages to host various social events (Park et al., 1925). In recent years, these spaces have also become key zones for the growing trend of micromobility (Mitchell et al., 2010; Abduljabbar et al., 2021; Oeschger et al., 2020), a term that refers to small, lightweight vehicles like electric scooters, e-bikes, and other mobile machines designed for short-distance travel. Micromobility is becoming an increasingly important solution for improving urban transportation efficiency, reducing environmental impact, and offering flexible alternatives to car ownership in cities.
@@
As shown in Figure 2 (Top), food delivery bots navigate on the sidewalk to accomplish the last-mile food delivery task, while elders and physically disabled people maneuver electronic wheelchairs and mobility scooters on the street. Various legged robots like robot dog Spot from Boston Dynamics and humanoid robot Optimus from Tesla are also forthcoming. We can imagine a future of such automated micromobility that harnesses advanced AI models to improve situational awareness and maneuver various mobile machines more intelligently and safely in complex urban environments.
@@
Simulation platforms have played a crucial role in enabling systematic and scalable training of embodied AI agents and safety evaluation before real-world deployment. However, most of the existing simulators focus either on indoor household environments (Puig et al., 2018; Kolve et al., 2017; Savva et al., 2019; Shen et al., 2021; Li et al., 2024; Gan et al., 2021) or outdoor driving environments (Krajzewicz et al., 2002; Li et al., 2022b; Dosovitskiy et al., 2017). For example, platforms like AI2-THOR (Kolve et al., 2017), Habitat (Savva et al., 2019), and iGibson (Shen et al., 2021) are designed for household assistant robots in which the environments are mainly apartments or houses with furniture and appliances; platforms like SUMO (Krajzewicz et al., 2002), CARLA (Dosovitskiy et al., 2017), and MetaDrive (Li et al., 2022b) are designed for research on autonomous driving and transportation focusing on roadways and highways. Yet, simulating complex urban environments for micromobility tasks, with diverse layouts, terrains, obstacles, and complex dynamics of pedestrians, is much less explored.
Figure 2: Motivation. (Top) Emerging automated micromobility. (Bottom) Unique challenges in micromobility.
@@
Distinct from the indoor household and driving tasks, micromobility plays an essential role in providing accessibility (e.g., electric wheelchairs) and convenience (e.g., food delivery bots) in the public urban space, while it also brings unique challenges for mobile machines and the underlying embodied AI agents. Let’s follow the adventure of a last-mile delivery bot, who aims to deliver a lunch order from a nearby pizzeria to the campus (Figure 2 (Bottom)). First, it faces a long-horizon task in large scale scenes across several street blocks, which span a significantly larger space than the indoor household environment. Second, it needs to deal with multifarious terrains, such as fragmented curbs and rugged ground caused by tree roots on sidewalks, which are seldom seen in indoor and driving environments. Then, it must safely navigate the cluttered street full of diverse obstacles like trash bins, parked scooters, and potted plants, which is absent in driving scenarios while with large obstacles variations compared to indoor scenes. In addition, it needs to handle dense pedestrians on sidewalks to avoid collisions, especially taking care of disabled people in wheelchairs, which do not exist in indoor and driving environments. Thus, the large scales, multifarious terrains, diverse obstacles, and dense pedestrians bring unique challenges to AI-driven mobile machines moving in cities, as well as the design of simulation environments for the training and evaluation of the embodied AI models.
@@
In this work, we present MetaUrban – a compositional simulation platform aiming to facilitate the research of AI-driven micromobility. First, we introduce Hierarchical Layout Generation, a procedural generation approach that can generate infinite layouts hierarchically from street blocks to sidewalks, functional zones, and object locations. It can generate scenes at an arbitrary scale with various connections and divisions of street blocks, obstacle locations, and complex terrains. Then, we design the Scalable Obstacle Retrieval, an automatic pipeline for acquiring an arbitrary number of high-quality objects with real-world distribution, to fill the urban space. We first compute the object category distribution from worldwide urban scene data to form a description pool. Then, with the sampled descriptions from the pool, we design a VLM-based open-vocabulary searching schema, which can effectively retrieve objects from large-scale 3D asset repositories. These two modules are critical for improving the generalizability of trained agents.
Finally, we propose the Cohabitant Populating method to generate complex dynamics in urban spaces. We first tailor recent 3D digital human and motion datasets to get 1,100 rigged human models, each with 2,314 movements. Then, to form safety-critical scenarios, we integrate Vulnerable Road Users (VRUs) like bikers, skateboarders, and scooter riders. As the subjects of micromobility, we include various mobile machines – the delivery bot, electric wheelchair, mobility scooter, robot dog, and humanoid robot. Then, based on path planning algorithms, we can get complex trajectories among hundreds of environmental agents simultaneously with collision and deadlock avoidance. Also, enabled by MetaUrban’s flexible user interfaces (mouse, keyboard, joystick, and racing wheel), users can directly apply human-operated trajectories to agents, which provides an easy way to collect demonstration data for agent training. In addition, we impose a series of traffic rules to regulate all agent behaviors. It is critical for enhancing the safety of the mobile agents.
[…]
@@ 2 Related Work
Many simulation platforms have been developed for embodied AI research, depending on the target environments – such as indoor homes and offices, driving roadways and highways, and crowds in warehouses and squares. We compare representative ones with the proposed MetaUrban simulator.
@@
Platforms for indoor environments are mainly designed for household assistant robots, emphasizing the affordance, realism, and diversity of objects, as well as the interactivity of environments. VirtualHome (Puig et al., 2018) pivots towards simulating routine human activities at home. AI2-THOR (Kolve et al., 2017) and its extensions, such as ManipulaTHOR (Ehsani et al., 2021), RoboTHOR (Deitke et al., 2020), and ProcTHOR (Deitke et al., 2022b), focus on detailed agent-object interactions, dynamic object state changes, and procedural scene generation, alongside robust physics simulations. Habitat (Savva et al., 2019) offers environments reconstructed from 3D scans of real-world interiors. Its subsequent iterations, Habitat 2.0 (Szot et al., 2021) and Habitat 3.0 (Puig et al., 2023b), introduce interactable objects and deformable humanoid agents, respectively. iGibson (Shen et al., 2021) provides photorealistic environments. Its upgrades, Gibson 2.0 (Li et al., 2021), and OmniGibson (Li et al., 2024), focus on household tasks with object state changes and a realistic physics simulation of everyday activities, respectively. ThreeDWorld (Gan et al., 2021) targets real-world physics by integrating high-fidelity simulations of liquids and deformable objects. However, unlike MetaUrban, these simulators are focused on indoor environments with particular tasks like object rearrangement and manipulation.
@@
Platforms for driving environments are mainly designed for autonomous vehicle research and development. Simulators like GTA V (Martinez et al., 2017), Sim4CV (Müller et al., 2018), AIRSIM (Shah et al., 2018), CARLA (Dosovitskiy et al., 2017), and its extension SUMMIT (Cai et al., 2020) offer realistic environments that mimic the physical world’s detailed visuals, weather conditions, and day-to-night transitions. Other simulators enhance efficiency and extensibility at the expense of visual realism, such as Udacity (Team, b), DeepDrive (Team, a), Highway-env (Leurent, 2018), and DriverGym (Kothari et al., 2021). MetaDrive (Li et al., 2022b) trades off between visual quality and efficiency, offering a lightweight driving simulator that can support the research of generalizable RL algorithms for vehicles. Although some of the simulators (Martinez et al., 2017; Dosovitskiy et al., 2017) involve traffic participants other than vehicles, such as pedestrians and cyclists, all of them focus on vehicle-centric driving scenarios and neglect the stage for urban micromobility – public urban spaces like sidewalks and plazas.
@@
Other than indoor and driving environments, social navigation platforms emphasize the social compatibility of robots. Simulators like Crowd-Nav (Chen et al., 2019), Gym-Collision-Avoidance (Everett et al., 2018), and Social-Gym 2.0 (Sprague et al., 2023), model scenes and agents in 2D maps, focusing more on the development of path planning algorithms. Other simulators, such as HuNavSim (Pérez-Higueras et al., 2023), SEAN 2.0 (Tsoi et al., 2022), and SocNavBench (Biswas et al., 2022), upgrade the environment to 3D space and introduce human pedestrians to support the development of more complex algorithms. Social navigation platforms focus on crowd navigation, with oversimplified objects and surrounding environmental structures in the scenes, making them not applicable to complex urban micromobility tasks. In contrast, MetaUrban supports large-scale urban space simulation with real-world scenes (such as street facilities and terrains), providing significantly rich semantics and superior complexity of environments. In addition, MetaUrban supports the cross-machine evaluation of generalizability and safety with different mechanical structures. These features make it a unique choice for urban micromobility.
In summary, none of the recent simulation platforms have been constructed for urban micromobility. The proposed simulator MetaUrban is the first simulator designed for AI-driven urban micromobility research. It differs from previous simulators significantly in terms of complex scenes (with large scales and multifarious terrains), diverse obstacles, vibrant dynamics, different types of mobile machines like delivery bots, electric wheelchairs, and mobility scooters, and intricate interactions simulated. Please refer to Appendix B.6 for a detailed comparison table with existing simulators in the dimensions of scale, sensor, and features, where MetaUrban shows a remarkable superiority. We believe MetaUrban can provide a lot of new research opportunities for AI-driven urban micromobility.
@@ 3 MetaUrban Simulator
MetaUrban is a compositional simulation platform that can generate infinite training and evaluation environments for AI-driven urban micromobility. We propose a procedural generation pipeline, as the basis of MetaUrban, for constructing infinite interactive scenes with different specifications. As shown in Figure 3, MetaUrban uses a structured description script to create urban scenes. Based on the script information about street blocks, sidewalks, objects, agents, and more, it starts with the street block map, then plans the ground layout by dividing different function zones, then places static objects, and finally populates dynamic agents.
@@
This section highlights three key designs in the MetaUrban simulator to support exhibiting three unique characteristics of urban spaces respectively – complex scenes (with large scales and multifarious terrains), diverse obstacles, and vibrant dynamics. Section 3.1 introduces Hierarchical Layout Generation, which can infinitely generate diverse layouts with different functional zone divisions, object locations, and terrains that are essential for the generalizability to scene diversity of agents. Section 3.2 introduces Scalable Obstacle Retrieval, which harnesses worldwide urban scene data to obtain real-world object distributions in different places, and then builds large-scale, high-quality static objects set with VLM-enabled open-vocabulary searching. It is crucial for further enhancing the generalizability to obstacle diversity of agents. Section 3.3 introduces Cohabitant Populating, in which we leverage the advancements in digital humans to enrich the appearances, movements, and trajectories of pedestrians and vulnerable road users, as well as incorporate other agents to form a vivid cohabiting environment. It is critical for improving the safety of the mobile agents.
@@ 3.1 Hierarchical Layout Generation
The complexity of scene layouts, i.e., the connection and categories of blocks, the specifications of sidewalks and crosswalks, the placement of objects, as well as the terrains, is crucial for enhancing the generalizability of trained agents maneuvering in public spaces. In the hierarchical layout generation framework, we start with a ground plan that samples categories of street blocks and divides sidewalks into different functional zones. Then, we allocate various objects procedurally conditioned on functional zones. Finally, we implement a terrain generation system to synthesize various ground conditions. With the above procedures, we can get infinite urban scene layouts with different specifications of sizes, object locations, and terrains.
@@
As shown in Figure 4 (Left), according to the Global Street Design Guide (Initiative & of City Transportation Officials, 2016) provided by the Global Designing Cities Initiative, we divide the sidewalk into four functional zones – building zone, frontage zone, clear zone, and furnishing zone. Based on their different combinations of functional zones, we further construct 7 typical templates for sidewalks (Figure 4 (Right)). To form a sidewalk, we can first sample the layout from the templates and then assign proportions for different function zones. For crosswalks, we provide candidates at the start and the end of each roadway, which support specifying the needed crosswalks or sampling them by a density parameter.
Terrain generation. We develop a procedural terrain generator by connecting sampled terrain primitives similar to the method adopted in (Lee et al., 2024), which uses the Wave Function Collapse (WFC) method (Gumin, 2016) to ensure smooth transitions between neighboring terrain primitives. We defined five types of terrain primitives, including slops, steps, stairs, ramps, and rough at different heights. After the mesh was generated, textures with different friction coefficients were added to the terrain to simulate different materials of the ground. This method allows for the generation of a wide variety of terrain combinations, reflecting the complex environments that agents may encounter.
@@
Object placement. After determining the ground layout, we can place objects on the ground. We divide objects into three classes. 1) Standard infrastructure, such as poles, trees, and signs, are placed periodically along the road. 2) Non-standard infrastructure, such as buildings, bonsai, and trash bins, are placed randomly in the designated function zones. 3) Clutter, such as drink cans, bags, and bicycles, are placed randomly across all functional zones. We can get different street styles by specifying an object pool while getting different compactness by specifying a density parameter. Please refer to Appendix B.1 for more details.
3.2 Scalable Obstacle Retrieval